OpenStudy (genny7):

How would you solve this? |5 – 3x| – 3 = 2

hartnn (hartnn):

start by adding 3 on both sides

OpenStudy (genny7):

okay

hartnn (hartnn):

what did u get after adding 3 on both sides ?

OpenStudy (genny7):

i add 3 on -3 and on 2 i canceled out -3 were i added 2 and 3 i got 5?

hartnn (hartnn):

good. |5x-3| =5 now if |a| = b then a=b or a =-b

hartnn (hartnn):

so if i have |5x-3| =5 i would get 5x-3 =5 or 5x-3 = -5 makes sense ?

OpenStudy (genny7):

yes now i understand thank you

hartnn (hartnn):

good,can you solve these ? 5x-3 =5 or 5x-3 = -5

myininaya (myininaya):

i see the word if sorry

hartnn (hartnn):

no problemo :)

OpenStudy (genny7):

Yes i can solve both of those equations

hartnn (hartnn):

good :D

OpenStudy (genny7):

thanks so much :)

hartnn (hartnn):

welcome ^_^ you seem to be new here, so \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (genny7):

Yes I am new here :) Thanks :)